# Spectral Differential Network Analysis for High-Dimensional Time Series

### v1

@@ 4 Simulations
In this section, we evaluate the performance of the proposed estimator $\hat{\Delta}$ using simulation studies. All simulations use VAR$(1)$ processes, with $p=54$, as this allows us to compute the true spectral density using results from Sun et al. (2018). We control the sparsity of the difference in inverse spectral densities by using a block diagonal structure. In all settings, the transition matrix in condition 1 is the same as that in condition 2 except the last $3\times 3$ block is multiplied by $-1$. The first setting uses similar coefficients as in Sun et al. (2018). In the second and third simulation settings, the transition matrix consists of one $51\times 51$ block and one $3\times 3$ block. The larger block was generated with 60% and 95% sparsity in the second and third settings, respectively. All simulations were performed using $n=100,200,500,1000,2000$ observations for both conditions. Smoothed periodograms were computed from data for each condition using a smoothing window of $M_{n}=\lceil n^{2/3}\rceil$ and then converted to the real space to generate $\hat{\Sigma}_{l}$. All computations were performed on a MacBook with a 2.3 GHz Quad-Core Intel Core i7 processor and 16 GB of memory. More information on the simulation setup, optimization techniques, and tuning parameter selection can be found in Appendix A.4.
@@
As previously noted, our method is the first to directly estimate the difference in inverse spectral densities and no direct competitors or state-of-the-art methods exist. Therefore, we conceptualized two potential alternatives to our SDD method and compare all methods in our simulations. The first alternative method is the Naïve method, which estimates $\Delta$ by taking the difference after estimating individual inverse spectral densities separately, $\hat{\Delta}_{\mathrm{N}}=\hat{\Sigma}_{1}^{-1}-\hat{\Sigma}_{2}^{-1}$. For the Naïve method, the graphical LASSO (GLASSO, Friedman et al., 2008) was used to estimate a sparse inverse spectral density in each condition. Specifically we used the fast implementation from Sustik and Calderhead (2012) available in the glassoFast package in R (GPL ($\geq$ 3.0)). To induce sparsity in individual estimates and (potentially) their difference, GLASSO was tuned for each condition separately using eBIC with $\gamma=0.5$ (Foygel and Drton, 2010). The next method directly thresholds the difference to induce sparse estimates. Specifically, we estimate the inverse spectral density in each condition and take the difference and then use Hard thresholding. The thresholds were tuned via BIC. Similar to Deb et al. (2024), all methods are compared based on the following metrics: Accuracy, Precision, Recall, and Relative Root Mean Square Error (RRMSE). The full definitions can be found in Appendix A.4.
@@
Simulation results for Accuracy and RRMSE in simulations 1 and 2 are reported in Figures 2 and 3. Results for Precision and Recall for simulations 1 and 2 can be found in Appendix A.4 Figures 8 and 9. Results from simulation 3 are in Appendix A.4, Figures 10 and 11. The full set of results for each of the simulations are in Appendix A.4 Tables 1, 2, 3. In general, SDD has better accuracy than the Naïve method with the difference in accuracy increasing as the sample size increases. This is likely due to the fact that, in general, SDD estimates much fewer edges than the Naïve method. This also results in fewer false positive edges but more false negatives, resulting in better precision for SDD at the expense of reduced recall; see Table 1 in Appendix A.4. Across all settings we also see that SDD generally has lower RRMSE compared to the Naïve method especially for large sample sizes. Overall, compared to the Naïve method, SDD identifies a similar number of true edges using a much smaller edge set. While Hard thresholding has slightly better accuracy than SDD in simulation setting 1 and better precision, SDD has much better RRMSE and recall across all settings and sample sizes. The standard errors of all methods are large due to the large variation across frequencies. To further support our analysis, we have also summarized the difference between SDD and each of the competing methods for each metric in Appendix A.4, Tables 456789.
@@ 5 Application to EEG Data
We apply our direct difference estimator (SSD) and competing estimators to electroencephalograms (EEG) data from Hatlestad-Hall et al. (2022). Data were recorded with a 64 channel EEG array for 111 healthy subjects at a sampling frequency of 1024 Hz. Four minutes of brain activity was recorded while subjects were resting with their eyes closed. For 42 subjects, a second session was recorded 2–3 months after the initial session. For our analysis, we used the pre-cleaned data provided in OpenNeuro Dataset ds003775 (Markiewicz et al., 2021). Specific cleaning steps can be found in Hatlestad-Hall et al. (2022). We also downsampled the data to 512 Hz.
@@
We compare the sparsity of the estimated differential network for SDD, the Naïve, and Hard threshold methods across the commonly-used Theta, Beta, Gamma and High-Gamma bands. Sparsity was defined as the proportion of entries in the estimated difference that were non-zero. Five or six evenly spaced frequencies were considered for each band. They were (in Hz): Theta: (4,5,6,7,8), Beta: (12, 16, 20, 24, 28), Gamma: (30, 40, 50, 60, 70), High-gamma: (80, 95, 110, 125, 140, 150). Results averaged over subjects and frequencies within each band are presented in Figures 4 and  5.
@@
6 Application to $\boldsymbol{\mu}\mathrm{\textbf{ECoG}}$ DATA
@@
We next apply our SDD estimator to optogenetic stimulation micro-electrocorticography ($\mu\mathrm{ECoG}$) data from Yazdan-Shahmorad et al. (2016, 2018); Bloch et al. (2022) [^1] The data is publicly available at https://datadryad.org/stash/dataset/doi:10.7272/Q61834NF. Studying how stimulation changes the brain network using such experiments can help design protocols to change the brain network in a targeted way. This can lead to treatments for conditions, such as schizophrenia or epilepsy, which are known to stem from abnormal brain connectivity networks (Bloch et al., 2022). A series of experimental and control sessions were conducted where neural activity from the surface of the brain of two adult male rhesus macaques was recorded. There were a total of 32 experimental and 4 control sessions. Each experimental session consisted of five 10-minute stimulation blocks (stimulation state) with a 5-minute recording block (resting state) before and after. Thus, each experimental session alternated between resting state and stimulation state blocks consisting of 6 total resting state blocks and 5 total stimulation state blocks. Control sessions had the same design as the experimental sessions except no stimulation occurred. Stimulation in experimental sessions was performed using optogenetics for simultaneous stimulation and recording. For each session, two locations corresponding to two electrodes were randomly chosen for optogenetic stimulation. To stimulate the brain, a 5ms laser pulse with a delay of either 100ms or 10ms was delivered every 200ms during the stimulation block. Recordings were obtained using a $92\mathrm{mm}^{2}$ micro-electrocorticography array of 96 electrodes at 24 kHz spanning the primary motor and sensorimotor (M1 and S1) regions. We downsampled to 500 Hz for our analysis.
@@
Prior work shows that stimulation induces changes across the entire brain connectivity network (Yazdan-Shahmorad et al., 2018; Bloch et al., 2022). We also study both how stimulation changes the connectivity between stimulation sites and across the entire network. However, instead of using a Naïve estimator of the connectivity changes as in Bloch et al. (2022) we use SDD. Using SDD, we estimate the change in inverse spectral densities from resting state to the following resting state (RS to RS) and from resting state to the following stimulation state (RS to SS). For each session there are 5 RS to RS differences (RS1 to RS2, RS2 to RS3, …, RS5 to RS6) and 5 RS to SS difference (RS1 to SS1, …, RS5 to SS5). Note that an estimated SDD coefficient of zero indicates no change while non-zero coefficients indicate a change. This could be either a change in magnitude of the connection, a new connection, or a lost connection. The frequencies we study in each band are the same as in Section 5. Results are averaged over all sessions, frequencies within each band, delay (control, 10ms, or 100ms) and RS to RS or SS to SS.
@@
Figure 6: Probability of Change in Edges Between Stimulation Sites by Frequency Band and Delay. Vertical bars indicate standard errors.
@@
First, we study how stimulation changes the edges between stimulation sites. We compute the probability of edge change between the stimulation sites for each band, delay, and RS to RS or RS to SS. Results are displayed in Figure 6. Compared to the control sessions, stimulation with a 10ms delay is more likely to induce a change in connectivity between the stimulation sites across all frequency bands. After stimulation ends (RS to RS analysis), this connectivity change is attenuated, but still persists. This is in line with prior work which showed a change in functional connectivity between stimulation sites in the stimulation state which persisted into the following resting state (Bloch et al., 2022). Interestingly, stimulation with a 100ms was also shown to induce a change in connectivity relative to the control sessions despite being outside the effective range for inducing a connectivity change. This is likely a result of the interaction between stimulation and the underlying network which has been shown to mediate connectivity changes (Bloch et al., 2022).
Next, we study how stimulation affects the entire brain network. Specifically we focus on how the distance of electrodes from stimulation sites impacts edge changes. While we expect electrode pairs closer to stimulation sites to have a higher likelihood of an edge change, the effect on the rest of the network is unclear as the effect of stimulation is mediated by the existing connectivity network (Bloch et al., 2022). For every electrode, we compute the distance to closest stimulation site and, for every pair of electrodes, we average these distances to obtain the average distance to the closest stimulation site. We then study how the relative probability of an edge change between electrodes varies by the mean stimulation distance. Since electrodes that are closer together are more likely to have an edge change even during control sessions (see Figure 12 in Appendix A.5), we need to adjust for this baseline probability of edge change. The relative probability of edge change thus adjusts for the baseline probability of an edge change for two electrodes at a specific distance apart. It can be interpreted as the probability of an edge change above/below what we would expect for two electrodes at the same distance apart in a control session. In Figure 7, we apply a generalized additive model (GAM) smoother to more clearly see trends in the data.
Figure 7: Relative Probability of Change in Edge by Average Distance to Stimulation Site. Lines are generated by smoothing using a GAM. Shaded areas represent 95% confidence intervals. A value of 0 (black line) represents no difference in probability of edge change relative to control sessions.
For both the 100ms and 10ms delays, we see similar curve shapes in the RS to RS and RS to SS analyses, suggesting that network changes induced during stimulation persist to the resting state. Across the Beta, Gamma, and High-Gamma bands, we see an increase in probability of an edge change for pairs of electrodes close to simulation sites for both the 10ms and 100ms delay, which is in line with prior works (Bloch et al., 2022). For medium distances, we see near zero relative probability of edge changes for all delays and all bands except Theta. However, in the Beta and Gamma bands, for pairs of electrodes that are $\approx 10-12.5$mm away from a stimulation site, the 100ms delay appears to induce more edge changes while the 10ms delay appears to have the opposite effect inducing fewer edge changes relative to baseline. In the Theta band, the 100ms delay induces fewer edge changes relative to baseline as distance from stimulation site increases. Conversely, the 10ms delay results in fewer edge changes relative to baseline in the Theta band for distances from stimulation site $<10$mm but produces more changes as distances from the stimulation site increase past 10mm. This suggests that when designing a stimulation protocol to induce connectivity changes, depending on the intended effect at different electrodes, either a 10ms or 100ms stimulation could be used.
@@ 7 Discussion
The SDD estimator was compared to the Naïve method of estimating sparse inverse the spectral densities and taking their difference, as well as a method that involved estimating the inverse spectral densities, forming the difference and then Hard thresholding the difference. The three methods were compared across three different simulation settings, where the true difference in inverse spectral densities was sparse. Comparison metrics included precision, recall, accuracy, and relative root mean square error (RRMSE). When compared to the other methods, SDD performed better in accurately identifying edges and non-edges and the SDD estimates had a lower relative root mean square error.
The SDD estimator was applied to both EEG and optogenetic stimulation $\mu\mathrm{ECoG}$ data. Using the EEG data we further validated SDD by showing it estimated sparser differences between networks that were closer in time, in line with expected behavior. Lastly SDD was applied to $\mu\mathrm{ECoG}$ data with optogenetic stimulation to study how the connectivity of the brain network changes in response to stimulation. We found that different stimulation parameters affected the connectivity network differently. This difference was likely due to the fact that the underlying brain network mediates changes in response to stimulation. We also found that the effects of stimulation persisted beyond the stimulation state.
Appendix A Appendix
@@
A.1 Additional Details for Methods
A.1.1 D-trace Loss
@@
where $\langle A,B\rangle=\mathrm{Tr}(AB^{T})$. Similar to Yuan et al. (2017), we can take the derivative with respect to $\Delta$ to see that $\Sigma_{1}^{-1}-\Sigma_{2}^{-1}$ minimizes the D-trace loss. More specifically, let
@@
Evaluating $\frac{\partial{L_{D}}}{\partial{\Delta}}$ at ${\Delta=\Sigma_{1}^{-1}-\Sigma_{2}^{-1}}$ yields 00, establishing that $\Delta=\Sigma_{1}^{-1}-\Sigma_{2}^{-1}$ minimizes $L_{D}$. Furthermore, the Hessian of $L_{D}$ with respect to $\Delta$ is $\frac{\partial^{2}{L_{D}}}{\partial{\Delta^{2}}}=(\Sigma_{1}\otimes\Sigma_{2}+\Sigma_{2}\otimes\Sigma_{1})/2$ where $\otimes$ is the kronecker product. Both $\Sigma_{1}$ and $\Sigma_{2}$ are positive definite, and thus $L_{D}$ is convex in $\Delta$ and $\Delta=\Sigma_{1}^{-1}-\Sigma_{2}^{-1}$ is the unique minimizer.
@@
A.2 Functional Dependence Framework

### v2

@@ 4 SIMULATIONS
In this section, we evaluate the performance of the proposed estimator $\hat{\Delta}$ using simulation studies. The code to reproduce these results as well as those in Section 5 are available at https://github.com/mikehellstern/spectral-differential-networks. All simulations use VAR$(1)$ processes, with $p=54$, as this allows us to compute the true spectral density using results from Sun et al. (2018). We control the sparsity of the difference in inverse spectral densities by using a block diagonal structure. In all settings, the transition matrix in condition 1 is the same as that in condition 2 except the last $3\times 3$ block is multiplied by $-1$. The first setting uses similar coefficients as in Sun et al. (2018). In the second and third simulation settings, the transition matrix consists of one $51\times 51$ block and one $3\times 3$ block. The larger block was generated with 60% and 95% sparsity in the second and third settings, respectively. All simulations were performed using $n=100,200,500,1000,2000$ observations for both conditions. Smoothed periodograms were computed from data for each condition using a smoothing window of $M_{n}=\lceil n^{2/3}\rceil$ and then converted to the real space to generate $\hat{\Sigma}_{l}$. All computations were performed on a MacBook with a 2.3 GHz Quad-Core Intel Core i7 processor and 16 GB of memory. More information on the simulation setup, optimization techniques, and tuning parameter selection can be found in Appendix D.
@@
As previously noted, our method is the first to directly estimate the difference in inverse spectral densities and no direct competitors or state-of-the-art methods exist. However, in this section we will discuss and compare several alternative methods. While not focused on directly estimating the difference, a related method is the joint graphical lasso with a fusion penalty (FGL, Danaher et al., 2014). This method aims to simultaneously estimate inverses in each condition that are believed to share structural similarities. To encourage similarities in the inverse estimates, FGL uses an $\ell_{1}$ penalty on their difference. In simulations, FGL was tuned using AIC from Danaher et al. (2014). We have also conceptualized two additional methods to compare to SDD. The first is a Naïve method, which estimates $\Delta$ by taking the difference after estimating individual inverse spectral densities separately, $\hat{\Delta}_{\mathrm{N}}=\hat{\Sigma}_{1}^{-1}-\hat{\Sigma}_{2}^{-1}$. For the Naïve method, the graphical LASSO (GLASSO, Friedman et al., 2008) was used to estimate a sparse inverse spectral density in each condition. Specifically we used the fast implementation from Sustik and Calderhead (2012) available in the glassoFast package in R (GPL ($\geq$ 3.0)). To induce sparsity in individual estimates and (potentially) their difference, GLASSO was tuned for each condition separately using eBIC with $\gamma=0.5$ (Foygel and Drton, 2010). The next method directly thresholds the difference in inverse spectral densities to induce sparse estimates. Specifically, we estimate the inverse spectral density in each condition and take the difference and then use Hard thresholding. The thresholds were tuned using the same eBIC as SDD given in (7). Similar to Deb et al. (2024), all methods are compared based on the following metrics: Accuracy, Precision, Recall, and Relative Root Mean Square Error (RRMSE). The full definitions can be found in Appendix D.
@@
Simulation results for Accuracy and RRMSE in simulations 1 and 2 are reported in Figures 2 and 3. Results for Precision and Recall for simulations 1 and 2 can be found in Appendix D Figures 8 and 9. Results from simulation 3 are in Appendix D, Figures 10 and 11. The full set of results for each of the simulations are in Appendix D Tables 1, 2, 3. In general, SDD has better accuracy than both FGL and the Naïve method with the difference in accuracy increasing as the sample size increases. This is likely due to the fact that, in general, SDD estimates much fewer edges than either method. This also results in fewer false positive edges but more false negatives, resulting in better precision for SDD at the expense of reduced recall; see Table 1 in Appendix D. Across all settings and sample sizes we also see that SDD generally has either lower or similar RRMSE compared to other methods. Overall, compared to FGL and the Naïve method, SDD identifies a similar number of true edges and achieves similar or lower RRMSE using a much smaller edge set. Compared to Hard thresholding, SDD achieves much better RRMSE and recall across all settings and sample sizes except setting 2 for large $n$ where Hard thresholding slightly outperforms SDD in terms of RRMSE. To further support our analysis, we have also summarized the difference between SDD and each of the competing methods for each metric in Appendix D, Tables 456789101112.
@@ 5 APPLICATION TO EEG DATA
We apply our direct difference estimator (SDD) and competing estimators to electroencephalograms (EEG) data from Hatlestad-Hall et al. (2022). Data were recorded with a 64 channel EEG array for 111 healthy subjects at a sampling frequency of 1024 Hz. Four minutes of brain activity was recorded while subjects were resting with their eyes closed. For 42 subjects, a second session was recorded 2–3 months after the initial session. For our analysis, we used the pre-cleaned data provided in OpenNeuro Dataset ds003775 (Markiewicz et al., 2021). Specific cleaning steps can be found in Hatlestad-Hall et al. (2022). We also downsampled the data to 512 Hz.
@@
We compare the sparsity of the estimated differential network for SDD, FGL, the Naïve, and Hard threshold methods across the commonly-used Theta, Beta, Gamma and High-Gamma bands. Sparsity was defined as the proportion of entries in the estimated difference that were non-zero. Five or six evenly spaced frequencies were considered for each band. They were (in Hz): Theta: (4,5,6,7,8), Beta: (12, 16, 20, 24, 28), Gamma: (30, 40, 50, 60, 70), High-gamma: (80, 95, 110, 125, 140, 150). Results averaged over subjects and frequencies within each band are presented in Figures 4 and  5.
@@
6 APPLICATION TO SIMULATED $\boldsymbol{\mu}\mathrm{\textbf{ECoG}}$ DATA
@@
We next apply our SDD estimator to simulated micro-electrocorticography ($\mu\mathrm{ECoG}$) data. The simulation design is based on the optogenetic stimulation $\mu\mathrm{ECoG}$ experiments from Yazdan-Shahmorad et al. (2016, 2018); Bloch et al. (2022). Studying how stimulation changes the brain network using such experiments can help design protocols to change the brain network in a targeted way. This can lead to treatments for conditions, such as schizophrenia or epilepsy, which are known to stem from abnormal brain connectivity networks (Bloch et al., 2022). Our simulations used a similar setup to Yazdan-Shahmorad et al. (2016). Specifically, a total of 32 experimental and 4 control sessions were simulated. Using the same session types as Bloch et al. (2022), of the 32 experimental sessions, 23 were generated using a simulated 10ms stimulation delay and 9 were generated with a 100ms delay. Each experimental session consisted of five 10-minute stimulation blocks (stimulation state) with a 5-minute recording block (resting state) before and after. Thus, each experimental session alternated between resting state and stimulation state blocks consisting of 6 total resting state blocks and 5 total stimulation state blocks. We simulated data using a sampling frequency of 500Hz.
Data in all sessions was simulated using a VAR(1) process with i.i.d standard normal errors. It is worth noting that due to the autoregressive nature of the process, autocovariance matrices between variables will not be diagonal in general despite using a diagonal covariance matrix for the errors (Lütkepohl, 2005). Since the experiments in Yazdan-Shahmorad et al. (2016) included a varying number of good electrodes out of the 96 total electrodes, we randomly selected between $p=56$ and $p=94$ good electrodes for each session. Coefficients in a base transition matrix ($A_{1}\in\mathbb{R}^{p\times p}$) were generated by randomly sampling from a standard normal distribution. To obtain a stable time series, the matrix $A_{1}$ was modified to have maximum eigenvalue $<1$. Similar to Yazdan-Shahmorad et al. (2016), two electrodes in each session were randomly chosen to be stimulated. The effect of stimulation was modeled by changing the coefficient in $A_{1}$ between simulation electrodes which is denoted as $A_{\mathrm{stim}}$. For control sessions, there was no stimulation so $A_{1}=A_{\mathrm{stim}}$. The effect of a 10ms and 100ms stimulation delay was simulated by changing the coefficient between stim sites to be $2\max(|A_{1}|)$ and $0.5\max(|A_{1}|)$, respectively. In this way, we simulate a weaker effect of the 100ms stimulation delay (Bloch et al., 2022). It is worth noting that, in contrast to the simulations in Section 4, the transition matrices are dense with no block structure. Thus, it is not immediately clear how the true spectral density changes as the transition matrices change.
During stimulation blocks, data were simulated using the corresponding $A_{\mathrm{stim}}$ matrix. To simulate a lasting, but weakened effect of stimulation on the following resting state, resting state blocks were simulated using a progressively increasing effect of stimulation. Specifically, resting state block 1 uses $A_{1}$ since no stimulation has occurred. The effect of stimulation for resting state blocks 2 through 6 linearly increases from $\approx 0.05$ to $\approx 0.6$ times the stimulation effect. For example, for a 10ms delay session, resting state block 2 was simulated using a VAR(1) process with transition matrix equal to $A_{1}$ where the coefficient between stimulation sites was replaced with $\approx 0.05*2\max(|A_{1}|)$.
@@
Using our simulated data, we study how stimulation changes the connectivity between stimulation sites. However, instead of using a naïve estimator of the connectivity changes as in Bloch et al. (2022), we use SDD. Using SDD, we estimate the change in inverse spectral densities from resting state to the following resting state (RS to RS) and from resting state to the following stimulation state (RS to SS). For each session there are 5 RS to RS differences (RS1 to RS2, RS2 to RS3, …, RS5 to RS6) and 5 RS to SS difference (RS1 to SS1, …, RS5 to SS5). Note that an estimated SDD coefficient of zero indicates no change while non-zero coefficients indicate a change. This could be either a change in magnitude of the connection, a new connection, or a lost connection. The frequencies we study in each band are the same as in Section 5. Results are averaged over all sessions, frequencies within each band, delay (control, 10ms, or 100ms) and RS to RS or SS to SS.
@@
Figure 6: RS to RS analysis. Probability of Change in Edges Between Stimulation Sites by Frequency Band and Delay. Vertical bars indicate standard errors.
Figure 7: RS to SS analysis. Probability of Change in Edges Between Stimulation Sites by Frequency Band and Delay. Vertical bars indicate standard errors.
@@
To study how stimulation changes the edges between stimulation sites, we compute the probability of edge change between the stimulation sites for each band, delay, and RS to RS or RS to SS. The RS to RS and RS to SS results are displayed in Figures 6 and 7 respectively. In the RS to SS analysis, stimulation with a 10ms or 100ms delay is more likely to induce a change in connectivity between the stimulation sites across all frequency bands compared to control sessions. After stimulation ends (RS to RS analysis), this connectivity change is attenuated, but still persists. This is in line with our simulation framework and prior work which showed a change in functional connectivity between stimulation sites in the stimulation state which persisted into the following resting state (Bloch et al., 2022).
@@ 7 DISCUSSION
The SDD estimator was compared to the joint graphical lasso with a fusion penalty (FGL, Danaher et al., 2014), the the Naïve method of estimating sparse inverse the spectral densities and taking their difference, and directly hard thresholding the difference. The four methods were compared across three different simulation settings, where the true difference in inverse spectral densities was sparse. Comparison metrics included precision, recall, accuracy, and relative root mean square error (RRMSE). When compared to the other methods, SDD performed better in accurately identifying edges and non-edges and the SDD estimates had a lower relative root mean square error. Using EEG data we further validated SDD by showing it estimated sparser differences between networks that were closer in time, in line with expected behavior.
8 CONCLUSION
@@
Acknowledgements
This work was partially supported by U.S. NIH grants R01-MH125429 and R01-GM133848.
Appendix
@@ Acknowledgements
Appendix A ADDITIONAL DETAILS FOR METHODS
A.1 D-trace Loss
@@
where $\langle A,B\rangle=\mathrm{Tr}(AB^{T})$. Similar to Yuan et al. (2017), we can take the derivative with respect to $\Delta$ to see that $\Sigma_{1}^{-1}-\Sigma_{2}^{-1}$ minimizes the D-trace loss.
More specifically, let
@@
Evaluating $\frac{\partial{L_{D}}}{\partial{\Delta}}$ at ${\Delta=\Sigma_{1}^{-1}-\Sigma_{2}^{-1}}$ yields 00, establishing that $\Delta=\Sigma_{1}^{-1}-\Sigma_{2}^{-1}$ minimizes $L_{D}$. Furthermore, the Hessian of $L_{D}$ with respect to $\Delta$ is $\frac{\partial^{2}{L_{D}}}{\partial{\Delta^{2}}}=(\Sigma_{1}\otimes\Sigma_{2}+\Sigma_{2}\otimes\Sigma_{1})/2$ where $\otimes$ is the kronecker product. Assuming both $\Sigma_{1}$ and $\Sigma_{2}$ are positive definite, $L_{D}$ is convex in $\Delta$ and $\Delta=\Sigma_{1}^{-1}-\Sigma_{2}^{-1}$ is the unique minimizer.
@@
Appendix B FUNCTIONAL DEPENDENCE FRAMEWORK